What should be added to both sides of the given equation to complete the square?
x^2-5x=4.

Answer by ewatrrr(24785)

 
Hi,
What should be added to both sides of the given equation to complete the square?
(5/2)^2 = 25/4 should be added to both sides
x^2-5x + 25/4 = 4 + 25/4
(x-5/2)(x-5/2) = 4 + 25/4
(x-5/2)^2 = 41/4